A sample of gas at 300 torr inside a steel tank is cooled from 500 °C to 0 °C. What is the final pressure of the gas in the steel tank? (round to one sig fig)


1
Expert's answer
2021-08-20T01:06:24-0400

P1/T1=P2/T2

T1 = 500 + 273 = 773K

T2 = 0 + 273 = 273K

300/773 = P2/273

P2 = 106 torr